A solution contains naphthalene (C10H8) dissolved in hexane (C6H14) at a concentration of 11.17 % naphthalene by mass.

Calculate the vapor pressure at 25 ∘C of hexane above the solution. The vapor pressure of pure hexane at 25 ∘C is 151 torr.

Express the answer using three significant figures.
Answer #1

Mass of naphthalene,m1 = 11.17 g

Molar mass of naphthalene,M1 = 128.17052 g/mol

Moles of napthalene,n1

n1 = m1/M1

n1= ((11.17 g) / (128.17052 g/mol))

n1 = 0.0866 moles

Mass of hexane,m2 = 88.83 g

Molar mass of hexane,M2 = 86.18 g/mol

Moles of hexane,n2

n2 = ((88.83 g )/(86.18 g/mol))

n2 = 1.03 moles

Mole fraction hexane = (1.03 / (1.03 + 0.0866))= 0.922

Vapor pressure of hexane = Pure vapor pressure x mole fraction

Vapor pressure of hexane  = 0.922×151 = 139 torr